OpenStudy (anonymous):

Someone check my work 1.A plant grew 1/8 inch last week and 3/8 inch this week. How many inches did the plant grow in the last two weeks? A. 1/8 inch B. 1/4 inch C. 1/2 inch ← the anwser I chose D. 5/8 inch 2. Sara had 7/8 pound of turkey. She used 2/3 pound for sandwiches. How many pounds of turkey does Sara have left A. 5/24 ← the anwser I chose B. 1/4 C. 5/12 D. 5/8 3. Jerome walked 2/3 mile in the morning and 11/12 mile in the evening. How many miles did he walk in all? A. 7/12 B. 19/24 C. 1 1/3 D. 1 7/12 ← the anwser I chose

OpenStudy (anonymous):

4.The length of a rectangle is 4/5 feet. The width is 5/8 feet. How much greater is the length of the rectangle than the width? A. 1/5 feet B. 7/40 feet ← the anwser I chose C. 3/20 feet D. 1/3 feet 5.A painter uses 2/3 gallon of paint on one job and 3/5 of the same paint on another. How many gallons of paint did the painter use for both jobs? A. 1/15 gallon B. 1 1/3 gallon C. 1 1/15 gallon D. 1 4/15 gallon ← the anwser I chose

OpenStudy (anonymous):

Can someone plz check my work??

OpenStudy (pixiedust1):

I'd be more than happy to check your work! What a wise student you are! :D

OpenStudy (anonymous):

Aww thank you!!

OpenStudy (pixiedust1):

1.) Correct! 3/8 + 1/8 = 4/8=1/2

OpenStudy (anonymous):

Ok great!

OpenStudy (the_fizicx99):

@PixieDust1 I'll let you explain xD, for 2 you have to find common denominators and subtract.

OpenStudy (pixiedust1):

Oh i'm sorry! If you wanna do it, go right ahead :)

OpenStudy (pixiedust1):

You may need to finish up though. i gtg in 6

OpenStudy (pixiedust1):

. 7/8 - 2/3 (to solve this , you must make both fractions have a common denominator. the lowest common denominator they can have is 24) (7x3)/(8x3) - (2x8)/(3x8) = 21/24 - 16/24 = 5/24 pounds

OpenStudy (pixiedust1):

So yup, you got #2 correct

OpenStudy (anonymous):

Thanks PixieDust1 and tHe_FiZiCx99 for taking time out of your busy evening to help me!

OpenStudy (pixiedust1):

2/3 + 11/12 (again, find the lowest common denominator before addition -> 12) (2*4)/(3*4) + (11*1)/(12*1) = 8/12 + 11/12 = 19/12 miles

OpenStudy (pixiedust1):

So yup, you got #3 correct!

OpenStudy (anonymous):

Yay!

OpenStudy (the_fizicx99):

Sure I'll explain question 2! An easy way to find common denominators is by multiplying them by each other. So 2/3 - 7/8 ... In order to find the CMD. Mulitply both the top and bottom of the first fraction by 8 so it should look like. 2*8 / 3*8 = 16/24 and 7*3/8*3 = 21/ 24 21/24 - 16/24 = 5/24

OpenStudy (ranga):

All of your answers are correct!

OpenStudy (pixiedust1):

because 19/12= 1 7/12

OpenStudy (pixiedust1):

Lol Ranga makes it all so simple lol

OpenStudy (anonymous):

lol :D

OpenStudy (anonymous):

Thanks to everyone!!! I really appericiate it!!

OpenStudy (the_fizicx99):

Yw

OpenStudy (ranga):

You are welcome.

OpenStudy (pixiedust1):

And I just checked the others over too! I gtg now, but they are all right :D tHe_FiZiCx99 can finish explaining if you need it :)

OpenStudy (anonymous):

Oh thanks alot! :D

OpenStudy (the_fizicx99):

ranga says they're all right, lol I trust his answers :)

OpenStudy (pixiedust1):

Not a problem! I'm always happy to help you :) You're so nice!

OpenStudy (ranga):

Nice and smart too!

OpenStudy (anonymous):

Awww thanks!

OpenStudy (anonymous):

Thanks PixieDust1 You are very nice and sweet